# A Comprehensive Analysis of the Effectiveness of Large Language Models as Automatic Dialogue Evaluators

### 24 Dec 2023 v1

@@ Abstract
Automatic evaluation is an integral aspect of dialogue system research. The traditional reference-based NLG metrics are generally found to be unsuitable for dialogue assessment. Consequently, recent studies have suggested various unique, reference-free neural metrics that better align with human evaluations. Notably among them, large language models (LLMs), particularly the instruction-tuned variants like ChatGPT, are shown to be promising substitutes for human judges. Yet, existing works on utilizing LLMs for automatic dialogue evaluation are limited in their scope in terms of the number of meta-evaluation datasets, mode of evaluation, coverage of LLMs, etc. Hence, it remains inconclusive how effective these LLMs are. To this end, we conduct a comprehensive study on the application of LLMs for automatic dialogue evaluation. Specifically, we analyze the multi-dimensional evaluation capability of 3022228 open-source models plus 2 proprietary models. recently emerged LLMs at both turn and dialogue levels, using a comprehensive set of 12 meta-evaluation datasets. Additionally, we probe the robustness of the LLMs in handling various adversarial perturbations at both turn and dialogue levels. Finally, we explore how model-level and dimension-level ensembles impact the evaluation performance. All resources are available at https://github.com/e0397123/comp-analysis.
@@ Introduction
The recent advancement in large language models (Brown et al. 2020; Chowdhery et al. 2022; Touvron et al. 2023a) coupled with refined alignment techniques (Wei et al. 2022a; Ouyang and et al. 2022) lead to an extensive array of general-purpose AI assistants that are capable of tackling a broad spectrum of NLP tasks. Harnessing their strong language understanding capability, LLMs, especially the family of instruction-tuned models, offer great promise as effective and generalized automatic dialogue evaluators. Several recent works (Huynh et al. 2023; Chen et al. 2023a; Liu et al. 2023; Lin and Chen 2023; Fu et al. 2023) report strong correlations of LLMs with human evaluation. Yet, the scope of their assessment is limited: (1) The coverage of LLMs is restricted, with a primary emphasis on proprietary models, such as OpenAI ChatGPT/GPT-4. Lately, there has been an exponential growth of open-source foundation models (Almazrouei et al. 2023; Touvron et al. 2023a; Scao et al. 2023) and ChatGPT-like LLMs (Taori et al. 2023; Chiang et al. 2023; Chen et al. 2023b), a timely survey is neccessary to examine their effectiveness as automatic dialogue evaluators. (2) The mode of evaluation primarily concentrates on correlation analysis with a limited number of meta-evaluation datasets. We not only conduct a comprehensive correlation analysis on 12 meta-evaluation datasets3336 turn-level and 6 dialogue-level respectively. along different evaluation dimensions [^4] Dimension refers to the quality aspect that is usually assessed in the human evaluation process, such as relevance at the turn level and dialogue coherence at the dialogue level., but also probe the robustness of LLMs against adversarial perturbations at both turn and dialogue levels.
@@
Turn-Level Datasets #Instance #Context Utt Doc Len IAA Range Reused Annotations Missing Annotations
Persona-USR (Mehri and Eskenazi 2020b) 300 9.3 98.4 / 12.0 $0.3\sim 0.7$ rel, int, und, ovr spe
Persona-Zhao (Zhao, Lala, and Kawahara 2020) 900 5.1 48.8 / 11.5 $>0.7$ ovr rel, int, und, spe
DailyDialog-Zhao (Zhao, Lala, and Kawahara 2020) 900 4.7 47.5 / 11.0 $>0.7$ rel, ovr int, und, spe
Topical-USR (Mehri and Eskenazi 2020b) 360 11.2 236.3 / 22.4 $0.5\sim 0.7$ rel, int, und, ovr spe
FED-Turn (Mehri and Eskenazi 2020a) 375 10.4 87.3 / 13.3 $0.5\sim 0.8$ rel, int, spe, und, ovr -
ConTurE-Turn (Ghazarian et al. 2022a) 1066 3.8 21.7 / 11.0 $\sim 0.3$ ovr rel, int, und, spe
Dialogue-Level Datasets #Instance #Utterances Doc Len IAA Range Reused Annotations Missing Annotations
IEval-Dial (Svikhnushina, Filippova, and Pu 2022) 500 6.0 74.4 - - coh, eng, inf, div, ovr
Persona-See (See et al. 2019) 500 12.0 91.2 - - coh, eng, inf, div, ovr
Reliable-Dial (Ji et al. 2022) 500 21.2 178.1 - - coh, eng, inf, div, ovr
ConTurE-Dial (Mehri et al. 2022b) 119 17.9 153.9 - - coh, eng, inf, div, ovr
FED-Dial (Mehri and Eskenazi 2020a) 125 12.7 116.8 $0.7\sim 0.8$ coh, eng, inf, div, ovr -
Human-Eval (Smith et al. 2022) 286 12.0 139.2 - - coh, eng, inf, div, ovr
Table 1: Details of the meta-evaluation datasets. The “Reused Annotations” columns indicate the dimensions with available human-annotated scores. The “Missing Annotations” column denotes the dimensions that need to be annotated by us. The doc length is the average #words per context/response for turn-level datasets and dialogue for dialogue-level datasets. The IAA range shows the range of inter-annotator agreements of available human annotations.
@@
### Dialogue: [[[[Here is the input dialogue for annotation]]]] ### Instruction: Rate the coherence, engagingness, diversity, informativeness, and overall quality of the input dialogue on a scale of 1 to 5 and just output the corresponding ratings. ### Output Format: coherence - x engagingness - x diversity - x informativeness - x overall - x ### Your Response: [[[[Here is GPT-4’s output]]]]
[…]
### Context: [[[[Here is the dialogue context]]]] ### Response: [[[[Here is the input response for annotation]]]] ### Instruction: Rate the context relevance, specificity, interestingness, understandability, and overall quality of the response on a scale of 1 to 5 and just output the corresponding ratings. ### Output Format: relevance - x specificity - x interestingness - x understandability - x overall - x ### Your Response: [[[[Here is GPT-4’s output]]]]
[…]
Table 2: The instruction template for prompting GPT-4 to annotate both dialogue-level and turn-level data instances. For our meta-evaluation of the proprietary models including ChatGPT and Palm-2 Bison, we also use this instruction template.
@@
Table 2 illustrates how we prompt GPT-4 to perform the annotation task. When calling the GPT-4 API, we set the temperature and top_p to 0.7 and 0.95 respectively. The annotation process is repeated five times to mimic the scenario of having multiple crowd workers annotate each data instance. The inter-annotator or inter-round agreement for GPT-4 is derived by averaging the pairwise Pearson correlations between the annotation scores from any two annotation rounds. In general, a good inter-annotator or inter-round agreement exceeding 0.65 is observed when annotating the missing dimensions of different datasets.
@@
### Dialogue: [[[[Here is the input dialogue for evaluation]]]] ### Instruction: Above is a conversation between a human ([H]) and a chatbot ([C]). Question: Are the responses of the chatbot informative? ### Your Answer: [[[[Here is LLM’s output in terms of “Yes” or “No”]]]]
[…]
### Context: [[[[Here is the dialogue context]]]] ### Response: [[[[Here is the input response for evaluation]]]] ### Instruction: Above is a dialogue context and the corresponding response. Question: Is the response relevant to the context? ### Your Answer: [[[[Here is LLM’s output in terms of “Yes” or “No”]]]]
[…]
Table 3: An instruction template example for prompting open-source LLMs to conduct dialogue-level and turn-level evaluation. (a) Evaluate the informativeness of the input dialogue and (b) Evaluate the contextual relevance of the input response.
@@
As we cannot obtain the output probabilities of the proprietary models at the time of paper preparation, we adopt the explicit scoring procedure to directly prompt them to produce multi-dimensional ratings of dialogues or responses. ChatGPT, Palm-2 Bison, and GPT-4 share the same instruction template outlined in Table 2. Due to their strong instruction-following abilities, we can easily extract the scores with matching heuristics. The rare erroneous cases are manually fixed. We repeat the scoring process of each proprietary model five times and apply the average score of the five runs as the corresponding $s^{dim}_{llm}$.
We do not apply the same approach for obtaining ratings from other open-source LLMs because their instruction-following abilities are weaker than the proprietary models. Sometimes, their generation becomes intractable. Instead, we follow Gupta et al. (2022) by applying an implicit scoring procedure. More specifically, given an instruction prompt input, we concentrate on the output probabilities related to the label words ”Yes” and ”No” as generated by the LLM. Then, we normalize the probability of “Yes” as $P(Yes)=P(Yes)/(P(Yes)+P(No))$ and $P(Yes)$ serves as the corresponding $s^{dim}_{llm}$. Greedy decoding is applied such that the generation process of the LLMs is deterministic. An example instruction template is shown in Table 3. The question design of each dimension is listed in Table 4. For different LLMs, we adapt the instruction template to match that used in their instruction-tuning process.
Dimension-Specific Question Design
Turn Level
Interestingness Question: Is the responses interesting?
Understandability Question: Is the responses understandable?
Relevance Question: Is the responses relevant to the context?
Specificity Question: Is the responses specific to the context?
Overall Question: Is the overall quality of the response good?
Dialogue Level
Engagingness Question: Is the conversation engaging?
Diversity Question: Are the responses of the chatbot diverse?
Table 4: Dimension-specific question design.
@@
Categories Models rel spe int und ovr avg coh eng div inf ovr avg
Vanilla LLaMA-7B (Touvron et al. 2023a) 0.114 0.056 0.073 0.113 0.018 0.075 0.341 0.142 0.049 0.166 0.109 0.161
LLaMA-13B (Touvron et al. 2023a) 0.442 0.240 0.333 0.348 0.402 0.353 0.522 0.425 0.237 0.324 0.404 0.382
LLaMA-2-7B (Touvron et al. 2023b) 0.241 0.187 0.178 0.134 0.259 0.200 0.225 0.227 0.044 0.261 0.292 0.210
LLaMA-2-13B (Touvron et al. 2023b) 0.303 0.122 0.161 0.269 0.225 0.216 0.318 0.362 -0.068 0.343 0.417 0.274
XGen-8K-7B (Nijkamp et al. 2023) 0.214 0.152 0.173 0.119 0.218 0.175 0.419 0.431 0.325 0.387 0.456 0.404
Falcon-7B (Almazrouei et al. 2023) 0.133 0.061 0.174 0.101 0.104 0.114 0.366 0.476 0.329 0.341 0.413 0.385
@@
OpenLLaMA-7B (Geng and Liu 2023) 0.158 0.008 0.105 0.154 0.184 0.122 0.402 0.452 0.270 0.336 0.429 0.378
OpenLLaMA-13B (Geng and Liu 2023) 0.231 0.112 0.175 0.184 0.226 0.186 0.425 0.506 0.239 0.342 0.471 0.397
Pythia-7B (Biderman et al. 2023) 0.145 0.097 0.083 0.106 0.146 0.115 0.079 0.204 0.026 0.094 0.220 0.124
BLOOM-7B (Scao et al. 2023) 0.082 0.021 0.124 0.099 0.122 0.089 0.224 0.271 0.137 0.217 0.277 0.225
Instruction-Tuned LLaMA-2-Chat-7B (Touvron et al. 2023b) 0.434 0.191 0.243 0.251 0.240 0.272 0.534 0.485 0.375 0.447 0.498 0.468
LLaMA-2-Chat-13B (Touvron et al. 2023b) 0.559 0.353 0.207 0.320 0.380 0.364 0.644 0.610 0.228 0.514 0.613 0.522
Alpaca-7B (Taori et al. 2023) 0.430 0.293 0.253 0.267 0.386 0.326 0.586 0.624 0.372 0.477 0.566 0.525
Vicuna-7B (Chiang et al. 2023) 0.368 0.096 0.221 0.100 0.219 0.201 0.490 0.468 0.279 0.488 0.482 0.441
Vicuna-13B (Chiang et al. 2023) 0.400 0.318 0.229 0.224 0.309 0.296 0.515 0.420 0.306 0.419 0.417 0.415
Falcon-Ins-7B (Almazrouei et al. 2023) 0.272 0.152 0.293 0.179 0.246 0.228 0.504 0.513 0.375 0.448 0.500 0.468
Tulu-13B (Wang et al. 2023) 0.585 0.427 0.369 0.350 0.488 0.444 0.659 0.661 0.326 0.518 0.657 0.564
Chimera-inst-chat-7B (Chen et al. 2023b) 0.489 0.276 0.373 0.309 0.368 0.363 0.563 0.599 0.439 0.525 0.607 0.547
Chimera-inst-chat-13B (Chen et al. 2023b) 0.547 0.449 0.377 0.366 0.404 0.428 0.582 0.671 0.432 0.585 0.563 0.567
Phoenix-inst-chat-7B (Chen et al. 2023b) 0.314 0.101 0.291 0.258 0.234 0.240 0.480 0.493 0.146 0.334 0.416 0.374
@@
XGen-8K-7B-Inst (Nijkamp et al. 2023) 0.272 0.293 0.267 0.145 0.265 0.248 0.502 0.515 0.308 0.417 0.506 0.450
ChatGLM-v2-6B (Zeng et al. 2023) 0.368 0.267 0.191 0.181 0.184 0.238 0.214 0.236 0.262 0.248 0.359 0.264
WizardLM-13B-V1.2 (Xu et al. 2023b) 0.463 0.422 0.390 0.245 0.314 0.367 0.572 0.580 0.455 0.513 0.632 0.550
Proprietary Palm-2 Bison (text-bison-001) 0.666 0.563 0.454 0.422 0.601 0.541 0.649 0.674 0.473 0.557 0.674 0.605
@@
It can be observed that ChatGPT and Palm-2 Bison are among the top 5 models across all the dimensions at both turn and dialogue levels. On average, they ranked first and second respectively. Out of the 32 open-source models available, Tulu-13B (Wang et al. 2023), Chimera-inst-chat-13B (Chen et al. 2023b), and Baize-v2-13B (Xu et al. 2023a) are distinguished as the top three performers at the turn-level. When it comes to dialogue-level performance, the top three models are Chimera-inst-chat-13B, Tulu-13B, and WizardLM-13B-V1.2 (Xu et al. 2023b). A significant gap can be observed between the proprietary models and the best open-source model. For example, at the turn level, ChatGPT outperforms Tulu-13B by 11% on average while at the dialogue level, ChatGPT outperforms Chimera-inst-chat-13B by 7.7%. The observations suggest the importance of the model scale and the quality of instruction data. The proprietary models are much larger than other open-source LLMs. They are trained on more sophisticated human-annotated instruction data while the open-source models are mainly instruction-tuned on data distilled from proprietary models.
@@
Models rel spe int und ovr avg
Baize-v2-13B 0.449 0.147 0.302 0.290 0.337 0.305
Tulu-13B 0.544 0.193 0.254 0.324 0.488 0.361
Chimera-inst-chat-13B 0.507 0.234 0.312 0.316 0.404 0.354
Palm-2 Bison 0.616 0.317 0.343 0.384 0.601 0.452
@@
Models coh eng div inf ovr avg
Tulu-13B 0.668 0.629 0.414 0.584 0.681 0.595
Chimera-inst-chat-13B 0.595 0.628 0.507 0.609 0.525 0.573
WizardLM-13B-V1.2 0.536 0.522 0.477 0.540 0.605 0.536
Palm-2 Bison 0.584 0.633 0.550 0.604 0.614 0.597
@@
Table 6: Dimension-specific Pearson correlation scores of GPT-4 vs other strong LLMs on human-annotated data. The best score in each dimension is highlighted in bold.
Performance of GPT-4
@@
Tulu-13B 0.492 0.488 0.656 0.657
Chimera-inst-chat-13B 0.476 0.404 0.680 0.563
Baize-v2-13B 0.417 0.337 0.578 0.469
WizardLM-13B-V1.2 0.408 0.314 0.598 0.632
Palm-2 Bison 0.602 0.601 0.693 0.674
@@
Table 8: Dimension-wise ensemble results of difference models for the overall quality evaluation.
@@
Perturbation Strategies Level Target Dims $\theta$ Descriptions
Random Response Turn Rel 0.3 The response is replaced with a response from other dialogues or the dialogue history.
Replace Pronoun Turn Rel 0.2 Pronoun usage in response is disturbed, such as swapping gender-specific pronouns with neutral ones.
Replace Named Entity Turn Rel 0.2 Named entities in response are replaced by other random entities.
Contradiction Turn Rel 0.2 The response’s meaning is altered to contradict the dialogue context.
Repetition Turn Und 0.2 A random portion of the response is re- peated multiple times to create make it unnatural.
Unnatural Paraphrase Turn Und 0.2 Non-stop English words are randomly sampled and replaced with unnatural synonyms from Wordnet.
Dullness Turn Int/Spe 0.2 The original response is replaced with a generic response or one of the speaker’s prior statements.
Order Shuffle Dialogue Coh 0.2 The order of utterances in a coherent dialogue is randomized.
Utterance Replacement Dialogue Coh 0.1 An utterance in a coherent dialogue is replaced with a random utterance drawn from other conversations.
Self-Contradiction Dialogue Eng 0.1 Human-crafted utterances are added to a dialogue such that one speaker contradicts their prior statements.
Utterance Duplication Dialogue Eng 0.1 The chatbot’s responses are modified by appending a segment of the response back to itself, making the responses unnatural.
Repeating Others Dialogue Eng 0.1 The chatbot’s responses are rephrased such that they consistently include portions of the human speaker’s previous statements.
Generic Utterance Dialogue Eng/Inf 0.2 ChatGPT is prompted to paraphrase chatbot replies in dialogue, such that the chatbot always gives generic responses.
Content Reduction Dialogue Inf 0.2 We prompt ChatGPT to simplify the chatbot responses in a dialogue such that the information density of the responses is low.
Table 10: The descriptions of different perturbation strategies. The “Level” column indicates whether the strategy is modifying a response or a multi-turn dialogue. The “Target Dims” column shows which dimensions the strategy targets. $\theta$ denotes the threshold value for each perturbation.
@@
Perturbation Strategies Level Data Source #Data Instances
Random Response Turn DailyDialog & PersonaChat 1894
Replace Pronoun Turn DailyDialog & PersonaChat 2919
Replace Named Entity Turn DailyDialog & PersonaChat 712
Contradiction Turn DailyDialog & PersonaChat 1094
Repetition Turn DailyDialog & PersonaChat 1993
Unnatural Paraphrase Turn DailyDialog & PersonaChat 369
Dullness Turn DailyDialog & PersonaChat 2811
Order Shuffle Dialogue FED-Dial 200
Utterance Replacement Dialogue FED-Dial 200
Self-Contradiction Dialogue DECODE 200
Utterance Duplication Dialogue FED-Dial 200
Repeating Others Dialogue FED-Dial 200
Generic Utterance Dialogue FED-Dial 200
Content Reduction Dialogue FED-Dial 200
Table 11: Adversarial Perturbation Data details
@@
Perturbations Palm-2 Bison ChatGPT Tulu-13B Chimera-13B Baize-13B WizardLM-13B L2-Chat-13B
@@
Table 12: Percentage of cases when the LLMs successfully detect a perturbation (Robustness Ratio $R$). The best ratio for each perturbation is highlighted in bold. L2-Chat-13B refers to LLaMA-2-Chat-13B.
@@
Motivated by prior works on applying perturbation strategies for metric robustness probing (Sai et al. 2021; Khalid and Lee 2022), we analyze the robustness of LLM evaluators with a series of adversarial strategies, which are presented in Table 10 and table 11 presents the data sources and statistics of our perturbation test suit.

### 20 Jan 2024 v2

@@ Abstract
Automatic evaluation is an integral aspect of dialogue system research. The traditional reference-based NLG metrics are generally found to be unsuitable for dialogue assessment. Consequently, recent studies have suggested various unique, reference-free neural metrics that better align with human evaluations. Notably among them, large language models (LLMs), particularly the instruction-tuned variants like ChatGPT, are shown to be promising substitutes for human judges. Yet, existing works on utilizing LLMs for automatic dialogue evaluation are limited in their scope in terms of the number of meta-evaluation datasets, mode of evaluation, coverage of LLMs, etc. Hence, it remains inconclusive how effective these LLMs are. To this end, we conduct a comprehensive study on the application of LLMs for automatic dialogue evaluation. Specifically, we analyze the multi-dimensional evaluation capability of 30 recently emerged LLMs at both turn and dialogue levels, using a comprehensive set of 12 meta-evaluation datasets. Additionally, we probe the robustness of the LLMs in handling various adversarial perturbations at both turn and dialogue levels. Finally, we explore how model-level and dimension-level ensembles impact the evaluation performance. All resources are available at https://github.com/e0397123/comp-analysis.
@@ Introduction
The recent advancement in large language models (Brown et al. 2020; Chowdhery et al. 2022; Touvron et al. 2023a) coupled with refined alignment techniques (Wei et al. 2022a; Ouyang and et al. 2022) lead to an extensive array of general-purpose AI assistants that are capable of tackling a broad spectrum of NLP tasks. Harnessing their strong language understanding capability, LLMs, especially the family of instruction-tuned models, offer great promise as effective and generalized automatic dialogue evaluators. Several recent works (Huynh et al. 2023; Chen et al. 2023a; Liu et al. 2023; Lin and Chen 2023; Fu et al. 2023) report strong correlations of LLMs with human evaluation. Yet, the scope of their assessment is limited: (1) The coverage of LLMs is restricted, with a primary emphasis on proprietary models, such as OpenAI ChatGPT/GPT-4. Lately, there has been an exponential growth of open-source foundation models (Almazrouei et al. 2023; Touvron et al. 2023a; Scao et al. 2023) and ChatGPT-like LLMs (Taori et al. 2023; Chiang et al. 2023; Chen et al. 2023b), a timely survey is neccessary to examine their effectiveness as automatic dialogue evaluators. (2) The mode of evaluation primarily concentrates on correlation analysis with a limited number of meta-evaluation datasets. We not only conduct a comprehensive correlation analysis on 12 meta-evaluation datasets2226 turn-level and 6 dialogue-level respectively. along different evaluation dimensions [^3] Dimension refers to the quality aspect that is usually assessed in the human evaluation process, such as relevance at the turn level and dialogue coherence at the dialogue level., but also probe the robustness of LLMs against adversarial perturbations at both turn and dialogue levels.
@@
Turn-Level Datasets #Data #Utt Doc Len IAA Range Reused Annotations Missing Annotations
Persona-USR (2020b) 300 9.3 98.4 / 12.0 $0.3\sim 0.7$ rel, int, und, ovr spe
Persona-Zhao (2020) 900 5.1 48.8 / 11.5 $>0.7$ ovr rel, int, und, spe
DailyDialog-Zhao (2020) 900 4.7 47.5 / 11.0 $>0.7$ rel, ovr int, und, spe
Topical-USR (2020b) 360 11.2 236.3 / 22.4 $0.5\sim 0.7$ rel, int, und, ovr spe
FED-Turn (2020a) 375 10.4 87.3 / 13.3 $0.5\sim 0.8$ rel, int, spe, und, ovr -
ConTurE-Turn (2022a) 1066 3.8 21.7 / 11.0 $\sim 0.3$ ovr rel, int, und, spe
Dialogue-Level Datasets #Data #Utt Doc Len IAA Range Reused Annotations Missing Annotations
IEval-Dial (2022) 500 6.0 74.4 - - coh, eng, inf, div, ovr
Persona-See (2019) 500 12.0 91.2 - - coh, eng, inf, div, ovr
Reliable-Dial (2022) 500 21.2 178.1 - - coh, eng, inf, div, ovr
ConTurE-Dial (2022b) 119 17.9 153.9 - - coh, eng, inf, div, ovr
FED-Dial (2020a) 125 12.7 116.8 $0.7\sim 0.8$ coh, eng, inf, div, ovr -
Human-Eval (2022) 286 12.0 139.2 - - coh, eng, inf, div, ovr
Table 1: Details of the meta-evaluation datasets. The “Reused” columns indicate the dimensions with available human-annotated scores. The “Missing” column denotes the dimensions that need to be annotated by us. The doc length is the average #words per context/response for turn-level datasets and dialogue for dialogue-level datasets. The IAA range shows the range of inter-annotator agreements of available human annotations. #Utt refers to the number of context utterances and dialogue utterances for the turn-level and dialogue-level datasets respectively.
@@
Figure 1: The instruction template for prompting GPT-4 to annotate both dialogue-level (top) and turn-level (bottom) data. For our meta-evaluation of the proprietary models including ChatGPT and Palm-2 Bison, we also use this instruction template.
@@
Figure 1 illustrates how we prompt GPT-4 to perform the annotation task. When calling the GPT-4 API, we set the temperature and top_p to 0.7 and 0.95 respectively. The annotation process is repeated five times to mimic the scenario of having multiple crowd workers annotate each data instance. The inter-annotator or inter-round agreement for GPT-4 is derived by averaging the pairwise Pearson correlations between the annotation scores from any two annotation rounds. In general, a good inter-annotator or inter-round agreement exceeding 0.65 is observed when annotating the missing dimensions of different datasets.
@@
Figure 2: An example for prompting open-source LLMs to evaluate the contextual relevance of the input response.
@@
As we cannot obtain the output probabilities of the proprietary models at the time of paper preparation, we adopt the explicit scoring procedure to directly prompt them to produce multi-dimensional ratings of dialogues or responses. ChatGPT, Palm-2 Bison, and GPT-4 share the same instruction template outlined in Figure 1. Due to their strong instruction-following abilities, we can easily extract the scores with matching heuristics. The rare erroneous cases are manually fixed. We repeat the scoring process of each proprietary model five times and apply the average score of the five runs as the corresponding $s^{dim}_{llm}$.
We do not apply the same approach for obtaining ratings from other open-source LLMs because their instruction-following abilities are weaker than the proprietary models. Sometimes, their generation becomes intractable. Instead, we follow Gupta et al. (2022) by applying an implicit scoring procedure. More specifically, given an instruction prompt input, we concentrate on the output probabilities related to the label words ”Yes” and ”No” as generated by the LLM. Then, we normalize the probability of “Yes” as $P(Yes)=P(Yes)/(P(Yes)+P(No))$ and $P(Yes)$ serves as the corresponding $s^{dim}_{llm}$. Greedy decoding is applied such that the generation process of the LLMs is deterministic. Figure 2 showcases an example to prompt open-source LLMs to evaluate response relevance [^8] All dimension-specific questions for prompting the open-source LLMs can be found in the Appendix.. For different LLMs, we adapt the instruction template to match that used in their instruction-tuning process.
@@
Models Rel Spe Int Und Ovr Avg Coh Eng Div Inf Ovr Avg
LLaMA-7B 0.114 0.056 0.073 0.113 0.018 0.075 0.341 0.142 0.049 0.166 0.109 0.161
LLaMA-13B 0.442 0.240 0.333 0.348 0.402 0.353 0.522 0.425 0.237 0.324 0.404 0.382
LLaMA-2-7B 0.241 0.187 0.178 0.134 0.259 0.200 0.225 0.227 0.044 0.261 0.292 0.210
LLaMA-2-13B 0.303 0.122 0.161 0.269 0.225 0.216 0.318 0.362 -0.068 0.343 0.417 0.274
XGen-8K-7B 0.214 0.152 0.173 0.119 0.218 0.175 0.419 0.431 0.325 0.387 0.456 0.404
Falcon-7B 0.133 0.061 0.174 0.101 0.104 0.114 0.366 0.476 0.329 0.341 0.413 0.385
@@
OpenLLaMA-7B 0.158 0.008 0.105 0.154 0.184 0.122 0.402 0.452 0.270 0.336 0.429 0.378
OpenLLaMA-13B 0.231 0.112 0.175 0.184 0.226 0.186 0.425 0.506 0.239 0.342 0.471 0.397
Pythia-7B 0.145 0.097 0.083 0.106 0.146 0.115 0.079 0.204 0.026 0.094 0.220 0.124
BLOOM-7B 0.082 0.021 0.124 0.099 0.122 0.089 0.224 0.271 0.137 0.217 0.277 0.225
LLaMA-2-Chat-7B 0.434 0.191 0.243 0.251 0.240 0.272 0.534 0.485 0.375 0.447 0.498 0.468
LLaMA-2-Chat-13B 0.559 0.353 0.207 0.320 0.380 0.364 0.644 0.610 0.228 0.514 0.613 0.522
Alpaca-7B 0.430 0.293 0.253 0.267 0.386 0.326 0.586 0.624 0.372 0.477 0.566 0.525
Vicuna-7B 0.368 0.096 0.221 0.100 0.219 0.201 0.490 0.468 0.279 0.488 0.482 0.441
Vicuna-13B 0.400 0.318 0.229 0.224 0.309 0.296 0.515 0.420 0.306 0.419 0.417 0.415
Falcon-Ins-7B 0.272 0.152 0.293 0.179 0.246 0.228 0.504 0.513 0.375 0.448 0.500 0.468
Tulu-13B 0.585 0.427 0.369 0.350 0.488 0.444 0.659 0.661 0.326 0.518 0.657 0.564
Chimera-7B 0.489 0.276 0.373 0.309 0.368 0.363 0.563 0.599 0.439 0.525 0.607 0.547
Chimera-13B 0.547 0.449 0.377 0.366 0.404 0.428 0.582 0.671 0.432 0.585 0.563 0.567
Phoenix-7B 0.314 0.101 0.291 0.258 0.234 0.240 0.480 0.493 0.146 0.334 0.416 0.374
@@
XGen-8K-7B-Inst 0.272 0.293 0.267 0.145 0.265 0.248 0.502 0.515 0.308 0.417 0.506 0.450
ChatGLM-v2-6B 0.368 0.267 0.191 0.181 0.184 0.238 0.214 0.236 0.262 0.248 0.359 0.264
WizardLM-13B-V1.2 0.463 0.422 0.390 0.245 0.314 0.367 0.572 0.580 0.455 0.513 0.632 0.550
Palm-2 (text-bison-001) 0.666 0.563 0.454 0.422 0.601 0.541 0.649 0.674 0.473 0.557 0.674 0.605
@@
It can be observed that ChatGPT and Palm-2 Bison are among the top 5 models across all the dimensions at both turn and dialogue levels. On average, they ranked first and second respectively. Out of the 28 open-source models available, Tulu-13B (Wang et al. 2023), Chimera-inst-chat-13B (Chen et al. 2023b), and Baize-v2-13B (Xu et al. 2023a) are distinguished as the top three performers at the turn-level. When it comes to dialogue-level performance, the top three models are Chimera-inst-chat-13B, Tulu-13B, and WizardLM-13B-V1.2 (Xu et al. 2023b). A significant gap can be observed between the proprietary models and the best open-source model. For example, at the turn level, ChatGPT outperforms Tulu-13B by 11% on average while at the dialogue level, ChatGPT outperforms Chimera-inst-chat-13B by 7.7%. The observations suggest the importance of the model scale and the quality of instruction data. The proprietary models are much larger than other open-source LLMs. They are trained on more sophisticated human-annotated instruction data while the open-source models are mainly instruction-tuned on data distilled from proprietary models.
@@
Rel Spe Int Und Ovr Avg
Baize 0.449 0.147 0.302 0.290 0.337 0.305
Tulu 0.544 0.193 0.254 0.324 0.488 0.361
Chimera 0.507 0.234 0.312 0.316 0.404 0.354
Palm-2 0.616 0.317 0.343 0.384 0.601 0.452
@@
Coh Eng Div Inf Ovr Avg
Tulu 0.668 0.629 0.414 0.584 0.681 0.595
Chimera 0.595 0.628 0.507 0.609 0.525 0.573
WizardLM 0.536 0.522 0.477 0.540 0.605 0.536
Palm-2 0.584 0.633 0.550 0.604 0.614 0.597
@@
Table 3: Dimension-specific Pearson correlation scores of GPT-4 vs other strong LLMs on human-annotated data. The best score in each dimension is highlighted in bold. All the open-source LLMs are the 13B variants.
@@
Tulu 0.492 0.488 0.656 0.657
Chimera 0.476 0.404 0.680 0.563
Baize 0.417 0.337 0.578 0.469
WizardLM 0.408 0.314 0.598 0.632
Palm-2 0.602 0.601 0.693 0.674
@@
Table 5: Dimension-wise ensemble results of difference models for the overall quality evaluation. All the open-source LLMs are the 13B variants.
@@
Level Source #Data Dims $\mathbf{\theta}$
RR Turn DD & PC 1894 Rel 0.3
RP Turn DD & PC 2919 Rel 0.2
RNE Turn DD & PC 712 Rel 0.2
Con Turn DD & PC 1094 Rel 0.2
Rep Turn DD & PC 1993 Und 0.2
UP Turn DD & PC 369 Und 0.2
Dul Turn DD & PC 2811 Int/Spe 0.2
OS Dial FED-Dial 200 Coh 0.2
UR Dial FED-Dial 200 Coh 0.1
SC Dial DECODE 200 Eng 0.1
UD Dial FED-Dial 200 Eng 0.1
RO Dial FED-Dial 200 Eng 0.1
GU Dial FED-Dial 200 Eng/Inf 0.2
CR Dial FED-Dial 200 Inf 0.2
Table 7: Adversarial Perturbation Data Statistics. DD & PC refer to DailyDialog & PersonaChat Respectively. RR, RP, RNE, Con, Rep, UP, Dul, OS, UR, SC, UD, RO, GU, and CR refer to Random Response, Replace Pronoun, Replace Named Entity, Contradiction, Repetition, Unnatural Paraphrase, Dullness, Order Shuffle, Utterance Replacement, Self-Contradiction, Utterance Duplication, Repeating Others, Generic Utterance, and Content Reduction respectively.
@@
Perturbations Palm-2 ChatGPT Tulu Chimera Baize WizardLM L2-Chat
@@
Table 8: Percentage of cases when the LLMs successfully detect a perturbation (Robustness Ratio $R$). The best ratio for each perturbation is highlighted in bold. All the open-source LLMs are the 13B variants and L2-Chat refers to LLaMA-2-Chat-13B.
@@
Motivated by prior works on applying perturbation strategies for metric robustness probing (Sai et al. 2021; Khalid and Lee 2022), we analyze the robustness of LLM evaluators with a series of adversarial strategies and table 7 presents the data sources and statistics of our perturbation test suit [^12] The detailed definitions of each perturbation strategy are outlined in the Appendix..